Please help me solve this equation, and take me through the steps . I have been told the answer but I am not sure how to do it on my own. 11/14x21/24
Answer by Alan3354(69443)

Please help me solve this equation, and take me through the steps . I have been told the answer but I am not sure how to do it on my own. 11/14x21/24
------------
There's no equal sign, so it's not an equation.
It's 2 fractions multiplied.
-------
11/14x21/24
= 11/14*7/8 = 11*7/(14*8)
= 77/112
= 11/16